Using kinetic friction to find force required to achieve a speed down a slope.

The coefficient of kinetic friction for a 22-kg bobsled on a track is 0.10. What force is required to push it down a 6.0 degree incline and achieve a speed of 60 km/hat the end of 75m?
Answer #1
Mass m = 22 kgThe coefficient of kinetic friction for a 22-kg bobsled on a track is μ = 0.10Angle θ = 6 oInitial speed u = 0Final speed v = 60 km/ h= 60 x 1000 m /3600 s= 16.666 m / sDistance S = 75 mFrom the relation v 2 - u 2 = 2aSAccleration a = [ v 2 - u 2] /2S= 1.851 m / s2Net force F =ma= 40.74 NRequired force F ' = F + frictional force= F + μmg cos θ= 40.74 + 21.44= 62.18 N
answered by: Emilee :(